i need to know how to solve word problems.. the price of a camera has been discounted 18%. the sale price is $340. find the list price.
Answer by likaaka(51)

Set your equation as so,
340/x=1-18%
340/x=82%, in other words 340 is 82% of what
Convert percentage to a decimal fraction
340/x=.82
340=.82x, solve for x
340/.82=x
x=414.6341repeated
Since your dealing in dollars, round to nearest hundredth
Original cost of camera is $414.63